OpenStudy (anonymous):

Given the function f(x) = 5(x+4) − 6, solve for the inverse function when x = 19. 1 68 72 84

OpenStudy (campbell_st):

do you know how to find the inverse...?

OpenStudy (campbell_st):

if you wait, someone always gives the answer. Good luck

OpenStudy (swissgirl):

Our first step would be to find the inverse The steps in finding the inverse is as follows: Replace all x with y's and all y's with x's Then simplify your equation And lastly isolate your y variable Voila you have your inverse

OpenStudy (swissgirl):

Our equation is as follows: f(x)=5(x+4)-6 which we can rewrite as y=5(x+4)-6 f(x) is just another way of saying y Firstly you would have to interchange the y and the x Can ya do that?